6. The world record for the men`s marathon is 2:03:38. If the distance is 42.195 km, what is the average velocity during the run?

6.

The world record for the men`s marathon is 2:03:38. If the distance is 42.195 km, what is the average velocity during the run?

Grade:12

3 Answers

Nirmal Singh.
10 years ago
total time = 2:03:38 hr = 2*3600+3*60+38 = 3818 sec
displacement = distance = 42.195 km = 42195 m
av. velocity = 42195/3818 = 11.0516 m/s
Thanks & regards,
Nirmal SIngh
Askiitians faculty
Anthony Cooper
6 years ago
THIS UTHER CID IS BAHD U Suk ?p ewhgargbvkrabhfkvarfbfab gvafb tdfhgwwbtacbbwtgbtrbrgbtgvbtgeywtrbfbwytb
Cecilia Caprino
3 years ago
Total time= 2:03:38 hr = 2*3600 + 3*60 + 38 = 7200+180+38= 7418 sec
displacement = distance = 42.195 km = 42195 m
average velocity = 42195/7418 = 5.688 m/s